OpenStudy (anonymous):

Is the simplified form of 2 to the square root 3 times the square root of 12 rational?

OpenStudy (anonymous):

Plz help me

OpenStudy (anonymous):

Rational ? Is the last part lol

OpenStudy (johnweldon1993):

\[\large 2^{\sqrt{3}} \times \sqrt{12}\] like that ?

OpenStudy (johnweldon1993):

Oh just is it rational...well lets find out

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@Compassionate @celestialdictator @emjaye108